# OpenSU3D: Open World 3D Scene Understanding using Foundation Models

### v1

@@ Abstract
In this paper, we present a novel, scalable approach for constructing open set, instance-level 3D scene representations, advancing open world understanding of 3D environments. Existing methods require pre-constructed 3D scenes and face scalability issues due to per-point feature vector learning, limiting their efficacy with complex queries. Our method overcomes these limitations by incrementally building instance-level 3D scene representations using 2D foundation models, efficiently aggregating instance-level details such as masks, feature vectors, names, and captions. We introduce fusion schemes for feature vectors to enhance their contextual knowledge and performance on complex queries. Additionally, we explore large language models for robust automatic annotation and spatial reasoning tasks. We evaluate our proposed approach on multiple scenes from ScanNet [1] and Replica [2] datasets demonstrating zero-shot generalization capabilities, exceeding current state-of-the-art methods in open world 3D scene understanding. Project page: https://opensu3d.github.io/
@@ I-A Motivation
Recent AI advancements have led to significant breakthroughs in open-set object detection and contextual understanding in 2D images, primarily due to pre-trained foundation models such as CLIP [3], SAM [4] and the integration of vision with language models [5, 6]. However, translating these successes to 3D scenes remains challenging. Current 3D approaches [7, 8, 9, 10, 11], while innovative, have not yet matched their 2D counterpart’s performance. Bridging this gap is crucial for applications requiring 3D interaction, such as robotics and digital twins, to revolutionize how we perceive and interact with the three-dimensional world.
I-B Limitations of Current Methods
@@
Recent works [7, 8, 9, 10, 11] have impressively integrated 2D foundation models for open-world 3D scene understanding. However, these methods have several limitations. Most of these methods are designed as batch-processing or non-incremental methods requiring the availability of complete 3D scene data before-hand, which is unrealistic in real world operation of many robotics applications. Primarily, these methods generate 3D feature vectors from 2D foundation models like CLIP [3, 12], but lack a universal strategy for 2D to 3D information extrapolation coming from other types of 2D foundation models. Furthermore, the creation of dense, per-point feature vectors not only poses memory and scalability issues but also complicates the critical task of isolating distinct entities within a scene, which is essential for practical applications. Most notably, the existing methods seem to work effectively for simple queries but lack the depth and contextual understanding required for more complex spatial queries/tasks.
I-C Proposed Approach
@@
We introduce a novel approach for constructing open-set 3D scene representation that addresses open vocabulary instance recall (object search), segmentation, annotation, and spatial reasoning. Our method leverages 2D foundation models to extract instance-level information from RGB-D images, using GroundedSAM [13] to obtain masks, bounding boxes, names, and prediction scores and GPT-4V [6] for detailed caption. For each instance, feature vectors are extracted from CLIP [3] at multiple scales and fused. Every instance in an image is assigned a unique ID and has an associated 2D segmentation mask. These masks are back-projected using depth and pose data to construct a per-image segmented 3D scene. Our approach tracks and updates the 2D and 3D segmentation masks and corresponding meta-information based on overlapping region based techniques, enabling efficient, scalable, and incremental 3D scene construction as the environment is explored. The instance-level representation and feature fusion schemes incorporate local context, aiding in distinguishing instances within the same class in relational queries.
I-D Key contributions
This study brings the following key contributions to the field of 3D scene understanding:
@@ I-D Key contributions
1. We introduce an incremental and scalable approach for open-set 3D scene understanding and instance segmentation that seamlessly integrates instance-level information from individual images, utilizing 2D foundational models, into a unified 3D scene representation.
2. We develop an innovative feature fusion formulation that enables the identification of instances within the same class through contextual queries and improves the overall open vocabulary 3D object search performance.
@@
3. We explore the use of large language models in conjunction with our constructed 3D scene representations for automatic annotation and complex spatial reasoning queries.
II RELATED WORK
II-A Foundation & Large Language Models
@@
Foundation models, trained on large scale data, have revolutionized AI by achieving unprecedented performance across diverse tasks. CLIP [3], BLIP [14], and similar models integrate visual and textual data to create unified representations, enabling improved performance in multimodal tasks such as image captioning, visual question answering, and cross-modal retrieval. In segmentation, LSeg [15], and OVSeg [16] offer promptable and open-vocabulary capabilities. Grounded models like GSAM [13] and SEEM [17], building on top of SAM [4], contextualize outputs by incorporating additional semantic and situational information, thereby enhancing the interpretative accuracy and relevance of model predictions. This allows for more precise and context-aware applications in tasks such as image segmentation, object detection, and scene understanding. Language models  [18, 19] have excelled in natural language tasks, while their integration with vision  [20, 6] advances open-world understanding. This work explores a generalizable approach for extracting and linking information between 2D images and 3D spaces, leveraging the capabilities of foundation and large language models.
II-B 3D Scene Segmentation
@@
Semantic segmentation is a key challenge in 3D vision. Methods like Voxblox++[21] and Kimera [22] merge semantics with SLAM, while others methods like Hydra[23], Scene Graph Fusion [24]) build 3D scene graph on top of semantically segmented scene, however, these techniques still operate within a closed-set paradigm. Recent works, including  [25] and SAM3D [26], identify 3D instances by analyzing overlapping semantically segmented points back-projected from RGB images and generates open-set fine-grained 3D masks non-incrementally. In contrast, [25] offers an incremental, closed-set sparse point map approach with fixed computation per update. We propose an incremental method using SAM’s 2D masks and region overlap-based techniques, generating fine-grained 3D instance masks with constant update computation. Additionally, our approach efficiently tracks 2D-3D mask IDs, facilitating efficient information transfer from per-image masks to full 3D scene.
II-C 3D Scene Understanding
3D scene understanding has evolved significantly by leveraging 2D vision-language models, mapping their rich features onto 3D spaces for open vocabulary queries. Initial works like OpenScene [8] and ConceptFusion [11] project 2D feature vectors from CLIPSeg [12] and CLIP [3] into 3D, demonstrating potential but posses computational and scalability issues due to dense, per-point representations. Recent work OpenMask3D [27] offers an instance-centric approach using Mask3D [28], avoiding scalability issues but it’s still non-incremental, requiring all the data beforehand. Additionally, some studies [11, 27] explore advanced feature engineering, fusing CLIP vectors from object-centric and larger image sections to capture both local and global context. Despite progress, challenges in computation, scalability, and non-incremental approaches necessitate more efficient, adaptable, and versatile solutions.
II-D 3D Spatial Reasoning
Global 3D spatial reasoning remains challenging in open-world scene understanding scenarios. Recent works 3DCLR [7], 3DLLM [9] and GroundedLLM [29] propose innovative approaches leveraging LLM for 3D reasoning, yet accurate 2D spatial reasoning is still challenging for models like GPT-4V [6], as shown in [30]. For 2D spatial reasoning, recent works like Set of Mark Prompting [30] explores prompting methods to directly use GPT4V [6]. We adapt this approach [30], employing LLMs for 3D spatial reasoning over our constructed scenes via strategic prompting.
II-E Concurrent Work
Concurrently, methods like Segment3D [31], OpenIns3D [32] and SayPlan [33] also tackle 3D scene understanding. OVSG [16] and ConceptGraph [34] closely align with our focus on incremental, scalable instance-based representation. However, we uniquely rely on geometric principles for merging 3D masks, and not CLIP similarity, resulting in better performance as shown in experiments. Additionally, instead of constructing separate scene graphs for spatial reasoning; we instead tap LLM’s innate reasoning abilities via tailored prompts over our constructed scene.
III METHOD
@@
Our method, processes a sequence of RGB-D images along with their poses to create an open set 3D scene representation for open world scene understanding tasks like, open vocabulary object retrieval, 3D segmentation, annotation and spatial reasoning. Illustrated in Fig. 1 the pipeline contains two main modules:
@@
2. 2D to 3D Fusion and Tracking: Creates a 3D semantic map from per-image 2D masks and associates 2D information into 3D space by tracking the corresponding IDs.
III-A Per-Image Feature Extraction
The feature extraction process, as shown in Fig. 2, begins with a sequence of RGB images, =$\mathcal{I}=\{I_{0},I_{1},I_{2},\ldots,I_{n}\}$. A subset =$\mathcal{I^{\prime}}=\{I_{0},I_{s},I_{2s},\ldots,I_{n}\}$ is sampled with a stride $s$ which ensures a reasonable overlap to minimize computational redundancy. For each image $I^{\prime}\in\mathcal{I^{\prime}}$, groundedSAM [13] is used to obtain 2D masks $M$, bounding boxes $BB$, and prediction scores $S_{\text{pred}}$. Crops of each instance, based on $bb\in BB$, are passed to GPT-4V [6] to get labels (names) $N$ and detailed captions $C$ describing the object.
Each instance is assigned a unique ID, and masks $M$ are updated to $M^{\prime}$ with these IDs and by adding a border of $px$ pixels around each mask to delineate entities. Feature vectors are extracted using the CLIP encoder in two stages:
@@ III-A Per-Image Feature Extraction
    A global feature vector $f_{\text{G}}$ is extracted for full image.
@@
    Instance-specific feature vectors $F=\{f_{\text{MS}}\}$ are created by first cropping the image at multiple scales based on scaling ratios $S_{r}=\{s_{r}\}_{k}$ and then fusing per-crop vectors $\{f\}_{k}$ with a multi-scale feature fusion scheme as discussed in Sec III-D.
@@
Figure 2: Feature extraction: Using Grounded-SAM, we extract masks and labels for each image. The labels are further refined and mask description is obtained using GPT-4V. Lastly, a global per-image CLIP feature and fused multi-scale per-mask CLIP features are extracted and stored.
III-B 2D to 3D Fusion & Tracking
As shown in Fig. 3, we initiate the fusion and tracking module by initializing an empty 3D point cloud for the complete 3D scene, represented as $\mathcal{P}_{\text{scene}}\in\mathbb{R}^{x,y,z,\text{ID}}$, and a global hash table $\mathcal{Q}$ for tracking the unique IDs, defined as: $\mathcal{Q}:\mathcal{Q}\mapsto\left\{\text{ID}\in\texttt{uniq}(\text{ID}\in\mathcal{P}_{\text{scene}}):\{\text{ID}\in\{M^{\prime}\}\}\right\}$
@@
$\mathcal{P}_{\text{frame}}=\left\{\left(T\cdot\left(D(u,v)\cdot K^{-1}\cdot\begin{pmatrix}u\\ v\\ 1\end{pmatrix}\right),M^{\prime}(u,v)\right)\right\}$ (1)
Using the bounds of $\mathcal{P}_{\text{frame}}$, we sample $\mathcal{P}^{\prime}_{\text{scene}}$ from $\mathcal{P}_{\text{scene}}$, containing only points within the bounds of $\mathcal{P}_{\text{frame}}$. A $KDTree$ search is performed, utilizing Euclidean distance function, $d(\cdot,\cdot)$ to matches points $\mathbf{p}$in $\mathcal{P}_{\text{frame}}$ with points $\mathbf{q}$in $\mathcal{P^{\prime}}_{\text{scene}}$. If $d(\mathbf{p},\mathbf{q})<\epsilon$, we group the indices corresponding to $\mathbf{p}\in\mathcal{P}_{\text{frame}}$ with $\mathbf{q}\in\mathcal{P}_{\text{scene}}$ to obtain the respective index pairs $\{(\mathbf{i}_{\text{frame}},\mathbf{i}_{\text{scene}})\}$ for all overlapping points. This search strategy limits the search space of $KDTree$ to only overlapping region, thereby requiring a constant computation (search space) per update.
To track and update matched IDs, similar to SAM3D’s [26] approach, we begin by obtaining a list of unique IDs $\{\text{ID}_{f}\}$ for each segment and a corresponding list denoting the total point count $\{c_{\mathcal{P}_{f}}\}$ of each segment of $\mathcal{P}_{\text{frame}}$.
For each segment in $\mathcal{P}_{\text{frame}}$ with $c_{\mathcal{P}_{f}}\in\{c_{\mathcal{P}_{f}}\}$, we utilize the index pairs $\{(\mathbf{i}_{\text{frame}},\mathbf{i}_{\text{scene}})\}$ to obtain the set of points from $\mathcal{P}_{\text{scene}}$ that overlaps with $\mathcal{P}_{\text{frame}}$. From these points, we derive a list of unique segment IDs $\{\text{ID}_{s}\}$ and their corresponding total point counts $\{c_{\mathcal{P}_{s}}\}$.
@@
If the overlap ratio satisfies a pre-defined threshold, i.e., $\text{OverlapRatio}\geq\rho$, we perform an ID replacement and update operation. Specifically, all $\text{ID}_{f}\in c_{\mathcal{P}_{f}}$ present in the $\mathcal{P}_{\text{frame}}$ are replaced with $\text{ID}_{s}\in\max(\{c_{\mathcal{P}_{s}}\})$ to get, $\mathcal{P^{\prime}}_{\text{frame}}$ which is then concatenated to $\mathcal{P}_{\text{scene}}$. Additionally, set of points from $\mathcal{P^{\prime}}_{\text{scene}}$ can also be deleted to retain constant sparsity, ensuring fixed computation requirement per update.
The updated IDs are then appended to the $\mathcal{Q}$, conversely, if the overlap ratio does not meet the threshold requirement, a new entry is added in $\mathcal{Q}$.
Figure 3: 2D-3D Fusion and Tracking: It tracks and updates IDs of each back-projected semantic mask associated with an image by efficiently assessing the overlap region with other semantic masks. Essentially, the tracked IDs are recorded and updated mask-projections are concatenated.
III-C Post Processing
The updated point cloud $\mathcal{P}_{\text{scene}}$ with updated corresponding IDs based on overalap ratio in $\mathcal{Q}$, along with per-image meta data is processed to construct an instance-centered 3D map of the processed whole scene $\mathcal{M}$, defined as:
=$\mathcal{M}=\left\{\left(\mathcal{P},n,c,f_{\text{MV}},bb_{3D},(x_{c},y_{c})\right)_{i}\middle|i\in\texttt{uniq}(\text{ID}\in\mathcal{P}_{\text{scene}})\right\}$ (3)
For each distinct 3D object $\mathcal{P}_{i}$, we perform DBSCAN based clustering to reduce noise and achieve fine-grained segmentation. Accordingly, 3D bounding boxes $bb_{3D,i}$ and their centroids $(x_{c},y_{c})_{i}$ are recomputed. For each multi-view image corresponding to the 3D object $\mathcal{P}{i}$, the 2D IDs $\mathcal{Q}[\text{ID}\in\mathcal{P}{i}]$, names $N^{\prime}$, captions $C^{\prime}$, prediction scores $S^{\prime}_{\text{pred}}$, and feature vectors $F^{\prime}$ are retrieved using $\mathcal{Q}$ for aggregation and fusion. The label $n_{i}$ and caption $c_{i}$ of each 3D instance $\mathcal{P}_{i}$ are assigned using the maximum prediction score $S^{\prime}_{\text{pred}}$ , alternatively the labels for each object-crops in images with top $m$ prediction scores $S^{\prime}_{\text{pred}}$ are refined using a LLM [6] with prompt: “assign a name to the object based on a given list of names”, yielding more accurate names $n^{\prime}_{i}$. Lastly, features vectors corresponding to multiple views $f_{\text{MV}}$ are obtained via multiview scheme as described in Sec. III-D by fusing feature vectors with top $m$ scores $S^{\prime}_{\text{pred}}$.
III-D Feature Fusion
Given a list of feature vectors $\{f\}_{k}$ from multiple scale crops of an instance in an image and feature vectors $\{f_{MS}\}_{m}$ corresponding to multiple view of images of a 3D instance, a simple and direct fusion scheme aggregates these feature vectors as shown below:

### v2

@@ Abstract
In this paper, we present a novel, scalable approach for constructing open set, instance-level 3D scene representations, advancing open world understanding of 3D environments. Existing methods require pre-constructed 3D scenes and face scalability issues due to per-point feature representation, additionally struggle with contextual queries. Our method overcomes these limitations by incrementally building instance-level 3D scene representations using 2D foundation models, and efficiently aggregating instance-level details such as masks, feature vectors, names, and captions. We introduce fusion schemes for feature vectors to enhance their contextual knowledge and performance on complex queries. Additionally, we explore large language models for robust automatic annotation and spatial reasoning tasks. We evaluate our proposed approach on multiple scenes from ScanNet [1] and Replica [2] datasets demonstrating zero-shot generalization capabilities, exceeding current state-of-the-art methods in open world 3D scene understanding. Project page: https://opensu3d.github.io/
@@ I INTRODUCTION
Recent advancements in AI, particularly in open-set object detection and contextual understanding of 2D images, owe much to pre-trained foundation models like CLIP [3], SAM [4], and the integration of vision with language models [5, 6]. However, extending these breakthroughs to 3D scenes remains a challenge. While innovative, current 3D methods [7, 8, 9, 10, 11] have not yet achieved the performance levels seen in 2D. Addressing this gap is critical for robotics, as it could transform how robots perceive, interact, and reason within the three-dimensional world.
@@
Recent efforts [7, 8, 9, 10, 11] have made strides in adapting 2D foundation models for open-world 3D scene understanding. However, they face key limitations. Many are designed for batch-processing or non-incremental tasks, requiring complete 3D scene data upfront—unrealistic in many real-world robotics applications. These approaches primarily derive per point 3D feature vectors from 2D models like CLIP [3, 12], but lack a generalized approach to extrapolate 2D information from other foundational models. Moreover, generating dense, per-point feature representations raises memory and scalability challenges while complicating the task of isolating distinct entities in a scene, critical for robotic operations. Most notably, the existing methods seem to work effectively for simple queries but lack the depth and contextual understanding required for more complex spatial queries and reasoning tasks.
I-A Overview of Contributions
@@
We present a novel approach for constructing open-set 3D scene representation addressing open vocabulary instance recall, segmentation, annotation, and spatial reasoning. Our method leverages 2D foundation models, specifically GroundedSAM [13] and GPT-4V [6], to extract instance-level information from RGB images. For each instance, CLIP [3] feature vectors are extracted and fused at multiple scales. Unique IDs are assigned to each instance in an image; corresponding 2D masks are back-projected using depth and pose data to construct a segmented 3D scene. Our approach tracks association between 2D and 3D masks by assessing overlapping region in 3D space, and updates 3D mask using corresponding meta-information from 2D images, enabling efficient, scalable, and incremental 3D scene construction. Additionally, feature fusion schemes incorporate local context, aiding in distinguishing instances within the same class for relational queries. This study brings the following key contributions to the field of 3D scene understanding:
@@
1. We introduce an incremental and scalable approach for open-set 3D scene understanding and instance segmentation, integrating instance-level information from 2D foundational models into a unified 3D representation.
2. We develop an innovative feature fusion formulation encompassing contextual information for improved instance identification through contextual queries.
@@
3. We explore the use of large language models in conjunction with our 3D scene representations for robust automatic annotation and complex spatial reasoning tasks
I-B Literature Review
I-B1 Foundation & Large Language Models
@@
Foundation models, trained on large-scale data, have transformed AI by achieving exceptional performance across varied tasks. CLIP  [3], and [14, 15] integrate visual and textual information into a unified representation, enabling multimodal tasks such as image captioning, visual question answering, and cross-modal retrieval. In segmentation,  [16, 17] provide promptable, open-vocabulary capabilities. Grounding approaches,  [13, 18], build on SAM [4] contextualize outputs by incorporating semantic and situational information, improving interpretative accuracy and relevance in tasks like image segmentation, captioning and object detection. Language models [19, 20] excel in natural language tasks, and their integration with vision [21, 6] further advances open-world comprehension. This work explores generalizable approach for extracting and linking information between 2D images and 3D spaces, leveraging the capabilities of foundation and large language models.
I-B2 3D Scene Segmentation
@@
Semantic segmentation remains a critical challenge in 3D robotics perception. While approaches like [22, 23, 24, 25] creates 3D metric semantic map, they are limited by closed-set paradigms. Recent works [26, 27] employ simplified approach, identifying 3D instances through overlapping 3D points from semantically segmented RGB images.  [27] produces non-incrementally open-set fine-grained 3D masks, while [26] creates an incremental, closed-set sparse 3D semantic map with fixed per-update computation. We propose an incremental method leveraging SAM’s [4] 2D masks and region overlap techniques to generate fine-grained 3D instance masks with constant per-update computation. Additionally, our approach efficiently tracks 2D-3D mask correspondence, facilitating effective instance-level information transfer from 2D foundation models to 3D scenes.
I-B3 Open Vocab. 3D Scene Understanding
Recent advancements in 3D scene understanding have leveraged 2D vision-language models to enable open-vocabulary interaction. While early works  [8, 11, 10, 28] demonstrated potential, they faced computational and scalability challenges due to dense feature representations. Subsequent approaches,  [29, 30], adopted instance-centric methods to address these issues. However, their non-incremental nature and reliance on pre-constructed 3D scene limits applicability in robotics. Additionally,  [11, 29] explores feature engineering to improve feature representation, fusing CLIP [3] vectors from object-centric and larger image sections but only shows performance on simple queries. Despite progress, challenges persist in computation, scalability, incremental processing, and contextual query handling, necessitating more efficient and adaptable solutions.
I-B4 3D Spatial Reasoning
Global 3D spatial reasoning remains a challenge in open-world scene understanding. While conventional scene graph approaches [24, 25] show limitations, recent works [7, 9, 31] leverage LLMs for 3D reasoning. However, even advanced models like GPT-4V [6] struggle with accurate 2D spatial reasoning [32]. Adapting the Set of Mark Prompting technique [32] for 3D environments, we employ LLMs for 3D spatial reasoning over constructed scenes through strategic prompting.
I-B5 Recent Related Works
Recent works [33, 34, 35, 30, 36, 37, 38] also addresses 3D scene understanding. Unlike  [36] and  [37], our method ensures semantic consistency in 3D masks through geometric overlap of 3D points rather than CLIP feature vectors, achieving better performance as shown in experiments. While  [35, 30], are non-incremental,  [33, 34] struggle with contextual queries.  [36, 37, 38] employ scene graphs and LLMs for spatial queries; in contrast, we explore in-context learning, leveraging large context length of LLMs and tailored prompting for complex spatial reasoning tasks.
II METHOD
@@
Our approach processes RGB-D image sequences with corresponding poses to generate an open-set 3D scene representation, supporting open-world tasks such as open-vocabulary object retrieval, 3D segmentation, annotation, and spatial reasoning. As shown in Fig.1, the pipeline consists of two key modules:
@@
2. 2D to 3D Fusion and Tracking: Creates a 3D semantic map from per-image 2D masks and associated 2D meta-information into 3D space by tracking the corresponding IDs.
II-A Per-Image Feature Extraction
The feature extraction process, begins with a sequence of RGB images, =$\mathcal{I}=\{I_{0},I_{1},I_{2},\ldots,I_{n}\}$. A subset $\mathcal{I^{\prime}}$ is sampled with a stride $s$ to minimize computational redundancy while ensuring a reasonable overlap. For each image $I^{\prime}\in\mathcal{I^{\prime}}$, groundedSAM [13] is used to obtain 2D masks $M$, bounding boxes $BB$, and prediction scores $S_{\text{pred}}$. Crops from image, based on $bb\in BB$, are passed to GPT-4V [6] with prompt “Give name and description to the object. Output e.g. ‘name’:‘description’” to get names (labels) $N$ and detailed captions $C$ describing the objects. Each instance is assigned a unique ID, and masks $M$ are updated to $M^{\prime}$ with these IDs as per-pixel labels, including a border of $px$ pixels around each mask to delineate entities. Feature vectors are extracted using the CLIP encoder in two stages:
@@ II-A Per-Image Feature Extraction
    A global feature vector $f_{\text{G}}$ is extracted from full image.
@@
    Instance-specific feature vectors $F=\{f_{\text{MS}}\}$ are created by first cropping the image at multiple scales based on scaling ratios $S_{r}=\{s_{r}\}_{k}$ and then fusing per-crop feature vectors $\{f\}_{k}$ with a multiscale feature fusion scheme as discussed in Sec. II-B2.
@@
Figure 2: Feature Extraction Module. For each instance in an image assigns a unique ID and extracts name, bounding box, detailed caption, prediction score, and CLIP [3] features.
II-B 2D to 3D Fusion & Tracking
We initiate the fusion and tracking module by initializing an empty 3D point cloud for the complete 3D scene, represented as $\mathcal{P}_{\text{scene}}\in\mathbb{R}^{x,y,z,\text{ID}}$, and a global hash table $\mathcal{Q}$ for tracking the unique IDs, defined as: $\mathcal{Q}:\mathcal{Q}\mapsto\left\{\text{ID}\in\texttt{uniq}(\text{ID}\in\mathcal{P}_{\text{scene}}):\{\text{ID}\in\{M^{\prime}\}\}\right\}$
@@
$\mathcal{P}_{\text{frame}}=\left\{T\left(D(u,v)\cdot K^{-1}\begin{pmatrix}u\\ v\\ 1\end{pmatrix}\right),M^{\prime}(u,v)\right\}$ (1)
Using the bounds of $\mathcal{P}_{\text{frame}}$, we sample $\mathcal{P}^{\prime}_{\text{scene}}$ from $\mathcal{P}_{\text{scene}}$, containing only points within the bounds of $\mathcal{P}_{\text{frame}}$. A $KDTree$ search is performed, utilizing Euclidean distance function, $d(\cdot,\cdot)$ to match points $\mathbf{p}$in $\mathcal{P}_{\text{frame}}$ with points $\mathbf{q}$in $\mathcal{P^{\prime}}_{\text{scene}}$. If $d(\mathbf{p},\mathbf{q})<\epsilon$, we group the indices corresponding to $\mathbf{p}\in\mathcal{P}_{\text{frame}}$ with $\mathbf{q}\in\mathcal{P}_{\text{scene}}$ to obtain the respective index pairs $\{(\mathbf{i}_{\text{frame}},\mathbf{i}_{\text{scene}})\}$ for all overlapping points. This search strategy limits the search space of $KDTree$ to only overlapping region, thereby requiring a constant computation (search space) per update.
To track and update matched IDs, similar to SAM3D’s [27] approach, we begin by obtaining a list of unique IDs $\{\text{ID}_{f}\}$ for each 3D segment and a corresponding list denoting the total point count $\{c_{\mathcal{P}_{f}}\}$ of each 3D segment in $\mathcal{P}_{\text{frame}}$.
For each 3D segment in $\mathcal{P}_{\text{frame}}$ with $c_{\mathcal{P}_{f}}\in\{c_{\mathcal{P}_{f}}\}$, we utilize the index pairs $\{(\mathbf{i}_{\text{frame}},\mathbf{i}_{\text{scene}})\}$ to obtain the set of points from $\mathcal{P}_{\text{scene}}$ that overlaps with $\mathcal{P}_{\text{frame}}$. From these points, we derive a list of unique segment IDs $\{\text{ID}_{s}\}$ and their corresponding total point counts $\{c_{\mathcal{P}_{s}}\}$.
@@
If the overlap ratio satisfies a pre-defined threshold, i.e., $\text{OverlapRatio}\geq\rho$, we perform an ID replacement and update operation. Specifically, all $\text{ID}_{f}\in c_{\mathcal{P}_{f}}$ present in the $\mathcal{P}_{\text{frame}}$ are replaced with $\text{ID}_{s}\in\max(\{c_{\mathcal{P}_{s}}\})$ to get, $\mathcal{P^{\prime}}_{\text{frame}}$ which is then concatenated to $\mathcal{P}_{\text{scene}}$. Additionally, set of points from $\mathcal{P^{\prime}}_{\text{scene}}$ can also be deleted to retain constant sparsity, ensuring fixed computation requirement per update. The updated IDs are then appended to the $\mathcal{Q}$; conversely, if the overlap ratio does not meet the threshold requirement, a new entry is added in $\mathcal{Q}$.
Figure 3: 2D-3D Fusion and Tracking. Tracks IDs of projected 3D semantic masks, associated with instance in image by assessing overlap in 3D space; tracked IDs are recorded and updated mask-projections concatenated.
II-B1 Post Processing
The point cloud $\mathcal{P}_{\text{scene}}$ with updated IDs, corresponding tracked overlapping IDs $\mathcal{Q}$, along with per-image metadata, is processed into instance-centric map
=$\mathcal{M}=\left\{\left(\mathcal{P},n,c,f_{\text{MV}},bb_{3D},(x_{c},y_{c})\right)_{i}\middle|i\in\texttt{uniq}(\text{ID}\in\mathcal{P}_{\text{scene}})\right\}$.
For each distinct 3D object $\mathcal{P}_{i}$, we perform DBSCAN based clustering to reduce noise and achieve fine-grained 3D masks. Accordingly, 3D bounding box $bb_{3D,i}$ and centroid $(x_{c},y_{c})_{i}$ are computed. For each multiview image corresponding to the 3D object $\mathcal{P}{i}$, names $N^{\prime}$, captions $C^{\prime}$, prediction scores $S^{\prime}_{\text{pred}}$, and feature vectors $F^{\prime}$ are retrieved using $\mathcal{Q}$ for aggregation and fusion. The label (name) $n_{i}\in N^{\prime}$ and caption $c_{i}\in C^{\prime}$, corresponding to the highest $S^{\prime}_{\text{pred}}$ are assigned to 3D instance $\mathcal{P}_{i}$. Alternatively, top $m$ names from $N^{\prime}$ based on prediction scores $S^{\prime}_{\text{pred}}$, are refined using a LLM [6] with prompt: “assign a single name to the object based on a given list of names”, yielding more accurate label (name) $n^{\prime}_{i}$. Lastly, feature vector corresponding to multiple views $f_{\text{MV}_{i}}$ is obtained via multiview feature fusion formulation (described in Sec. II-B2) by top $m$ fusing feature vectors with based on scores $S^{\prime}_{\text{pred}}$.
II-B2 Feature Fusion
Given a list of feature vectors $\{f\}_{k}$ from multiple scale crops of an instance in an image and feature vectors $\{f_{MS}\}_{m}$ corresponding to multiview image of a 3D instance, a simple and direct fusion scheme aggregates these feature vectors as shown below: